could someone please help me solve this equation?
(x-5)^2=6

Answer by checkley71(8403)

(X-5)^2=6 OR X^2-10X+25=6 OR X^2-10X+19 OR (X-7.449)(X-2.550)=0 OR
X=7.449 & X=2.550 (USING THE QUADRATIC EQUATION)